# MMTU: A Massive Multi-Task Table Understanding and Reasoning Benchmark

### v3

@@ 1 Introduction
Remarkable progress has been made in foundation models [foundationmodels, gpt3, r1, llm-llama, llm-palm-2], partly thanks to an expanding array of large-scale benchmarking efforts. Prominent examples include benchmarks for general language understanding (e.g., GLUE [glue], Super-GLUE [superglue], BIG-Bench [big-bench], MMLU [mmlu], MMLU-pro [mmlu-pro]), as well as benchmarks focused on coding and STEM reasoning (e.g., SWE-bench [benchmark-swebench], GPQA-diamond [benchmark-gpqa], AIME [benchmark-aime], LiveCodeBench [benchmark-livecodebench]). These efforts have played a critical role in deepening our understanding and accelerating the progress of foundation models.
Tables and table-based use cases are central to many real-world applications, including spreadsheets [excel, Google-sheets], databases [database-oracle, database-sql-server], and computational notebooks [jupyter, databricks-notebook], which often require expert-level users such as data engineers, analysts, and database administrators to operate. While LLMs have shown great promise in working with tables [excel-copilot, database-copilot, gemini-google-bigquery], existing evaluations of table-tasks remain narrow in scope – primarily focusing on NL-to-SQL [ns-bird, ns-archer, ns-spider] and Table-QA [tqa-wikiqa, tqa-hitab, tqa-tabfact] – and fail to reflect the broader spectrum of real-world tasks that professional users face.
@@
Our evaluation shows that the complex and technical nature of MMTU demands a combination of capabilities – including table understanding, reasoning, and coding – that remain challenging for today’s frontier models. Even the top performing models, such as OpenAI GPT-5 and DeepSeek-R1 , achieve only 69.6% and 57.9% on MMTU, suggesting substantial room for improvement, and highlighting MMTU as a strong testbed for models aspiring toward general human-level intelligence, e.g., to meet or surpass the top 10% of skilled adults in diverse technical tasks like those characterized in [agi-levels].
@@
  • Reasoning-focused models, such as OpenAI GPT-5 and DeepSeek R1, show a clear advantage over general-purpose chat models like GPT-5-Chat and DeepSeek-V3. The top reasoning models outperform the top chat models by over 10 percentage points (Table 3), underscoring the complexity of the tasks (which often require coding in SQL/Pandas) and the importance of reasoning in MMTU.
@@
Table Transform table-transform-by-relationalization (TTBR) Relationalize a table using transformation Acc [ttbr-auto-tables, transform-table-barowy2015flashrelate, transform-table-foofah, xiong2025tablecanoniser, autoprep] 230
table-transform-by-output-schema (TTBS) Synthesize transformation by output schema Acc [tbps-auto-pipeline, sharma2023automatic, sharma2025datamorpher] 685
table-transform-by-output-table (TTBT) Synthesize transformation by input/output tables Acc [ttbt-autopandas, ttbt-scythe, transform-table-qbo, table-transform-sql-by-output] 86
Table Matching Entity matching (EM) Match rows refer to the same semantic entity Acc [em-magellandata, em-deepmatcher, em-auto-em, em-ditto, em-book] 4228
Schema matching (SM) Match columns refer to the same concept F1 [sm-valentine, sm-smurf, sm-other-1, schema-matching-cupid] 669
Head value matching (HVM) Match column-headers with cell-values Acc [table-gpt, schema-mapping-survey] 900
Data Cleaning data-imputation (DI) Predict missing values in tables Acc [webtables, tablib, table-gpt] 1971
error-detection (ED) Detect erroneous cells in tables F1 [error-detection-survey, error-detect-uni-detect, error-detect-baran, error-detect-auto-test] 1891
list-to-table (L2T) Split lists of undelimited values into table Acc [l2t-tegra, list-extraction-2, list-extraction-google, list-other-1] 957
Table Join semantic-join (SJ) Predict semantic join between two tables Acc [sj-sema-join, sj-dataxformer, semantic-join-dtt, auto-map] 130
equi-join-detect (EJ) Predict equi-joins between a set of tables F1 [ej-auto-bi, ej-chen2014fast, ej-fk-ml, ej-hpi] 494
Column Transform program-transform-by-example (PTBE) Program transformation by input/output examples Acc [ptbe-tde, ptbe-transform-prose, transform-column-devlin2017robustfill, transform-column-singh2016blinkfill, he2018transform] 558
formula-by-context (FBC) Predict formula based on table context Acc [fbc-auto-formula, chen2021spreadsheetcoder] 3513
semantic-transform-by-example (STBE) Predict semantic transformations by examples Acc [sj-sema-join, sj-dataxformer, semantic-join-dtt] 131
Column Relationship arithmetic-relationship (AR) Predict arithmetic-relationship (AR) in tables F1 [explain-relationship, Auto-relate] 818
functional-relationship (FR) Predict functional-relationship (FR) in tables F1 [Auto-relate, fd-1, fd-2] 267
string-relationship (SR) Predict string-relationship (SR) in tables F1 [Auto-relate, ptbe-tde, transform-col-flashfill] 765
Table understanding Needle-in-a-haystack-table (NIHT) Retrieve cell content in a table Acc [tablib] 999
Needle-in-a-haystack-index (NIHI) Retrieve index based on cell value Acc [tablib] 1000
NL-2-code NL-2-SQL (NS) Translate natural-language into SQL Acc [ns-archer, ns-bird, ns-kaggledbqa, ns-spider, ns-wikisql] 2489
Table QA Table Question Answering (TQA) Answer questions based on tables Acc [tqa-tablebench, tqa-wikiqa, tqa-finqa, tqa-hitab] 1793
Fact Verification (FV) Verify facts based on tables Acc [tqa-tabfact, tablefact-2, tablefact-3] 916
KB Mapping Column type annotation (CTA) Predict KB types based on column content Acc [semtab19, cta-autotype, cta-doduo, cta-sato, cta-sherlock] 881
Column property annotation (CPA) Predict KB property for a pair of columns Acc [semtab19, cpa-2, cpa-3] 873
Cell entity annotation (CEA) Predict KB entity for a table cell Acc [semtab19, turl, cea-1, cea-2, cea-3] 892
@@ 2 Related Work
Large-scale benchmarks for foundation models. The rapid advancement of foundation models has made benchmark evaluation ever more important. Prominent large-scale benchmarks, such as GLUE [glue] (9 NLP tasks), Super-GLUE [superglue] (10 NLP tasks), Big-BENCH [big-bench] (204 tasks), MMMU [mmlu] (15,908 questions), MMMU-pro [mmlu-pro] (12,032 questions), MMLU [mmmu] (11,550 multi-modal questions) etc., offer comprehensive evaluations of model capabilities. However, as models improve rapidly, benchmarks can become saturated quickly (e.g., in the matter of a few years), prompting newer and more challenging benchmarks [superglue, mmlu-pro]. All of these benchmarking efforts have nevertheless played a crucial role in measuring and stimulating the development of foundation models.
@@
Benchmarks for reasoning. Reasoning has recently emerged as a key challenge for foundation models. Beyond general intelligence benchmarks (e.g., GPQA Diamond [benchmark-gpqa], AGIEval [benchmark-agieval], HLE [benchmark-hle]), there are specialized benchmarks targeting mathematical reasoning (e.g., AIME [benchmark-aime], MathVista [benchmark-mathvista], IMO [benchmark-imo]), coding (e.g., SWE-bench [benchmark-swebench], CodeForce [benchmark-codeforce], LiveCodeBench [benchmark-livecodebench], IOI [benchmark-ioi]), and multi-modal reasoning (e.g., MMMU [mmmu], ARC [benchmark-arc]). These benchmarks have become important tools for evaluating models’ ability to tackle complex reasoning tasks, but can also get saturated quickly (e.g., GSM8k [gsm8k], Math500 [math500], HuamEval [humaneval]), making it necessary to create new and more challenging benchmarks.
@@
Existing benchmarks relating to tables. Given the importance of table data, benchmarks have been developed in the ML and NLP community to evaluate model ability on tables, which however usually focus on a small set of table tasks such as NL-2-SQL [ns-archer, ns-bird, ns-kaggledbqa, ns-spider, ns-wikisql] and Table-QA [tqa-tabfact, tqa-tablebench, tqa-wikiqa, tqa-finqa, tqa-hitab]. In contrast, MMTU expands the scope of current evaluations by incorporating 19 additional table tasks drawn from decades of research in communities such as data management and programming languages, resulting in a more comprehensive evaluation framework for assessing LLM capabilities on tabular data.
More recently, spreadsheet-centric benchmarks have emerged, including Spreadsheet-Bench [ma2024spreadsheetbench], SheetCopilotBench [li2023sheetcopilot], and Sheet-RM [chen2024sheetagent]. While these are important in the spreadsheet domain, these efforts are typically limited in scale (containing a few hundred cases), and are closely tied to specific file formats (e.g., .xlsx) and software environments. In comparison, MMTU focuses on general-purpose tabular data that applies broadly across spreadsheet, database, and computational notebook settings, enabling more scalable and format-independent evaluation of foundation models.
@@ 3.2 Data curation workflow
Literature survey. To ensure our table tasks reflect real-world challenges, we draw on our experience working on related problems, that many challenging predictive table tasks have been studied in the decades worth of computer science research, particularly in data management (SIGMOD/VLDB), programming languages (PLDI/POPL), and web data (WWW/WSDM) communities. We conduct a systematic survey of publications from these venues over the past two decades, leveraging a combination of keyword searches of paper titles, and DeepResearch-like tools [deepresearch] (where we specify detailed requirements for papers in these venues), to arrive at a promising set of papers and possible candidate tasks.
@@
  • (1) Real user-facing tasks, involving data tables that would otherwise require expert-level humans to perform. (We therefore exclude system-level predictive table tasks focused on performance improvements, such as query optimization [qo-1, qo-2] and cardinality estimation [cardinality-1, cardinality-2]);
  • (2) Objectively evaluable tasks, that come with unique manually-labeled ground truth. (We therefore exclude tasks such as table summarization [table-summary-2, table-summary-3, table-summary-google, table-summary-4, table-summary-5] and table augmentation [table-augmentation-2, table-augmentation-infogather], which lack unique ground-truth and may require subjective fuzzy LLM-based evaluations);
@@
To ensure the quality of MMTU, we conduct an LLM-based quality check using o4-mini. The model is prompted to evaluate each question for two aspects: (1) whether the question is ambiguous (since tasks like NL-code can sometimes be inherently ambiguous [saparina2024ambrosia]), and (2) whether the ground-truth answer is correct. Approximately 8% of the original questions were flagged as either ambiguous or incorrect and subsequently removed to enhance benchmark quality.
@@
Evaluation framework. In contrast to benchmarks like MMLU [mmlu] and MMMU [mmmu], which primarily use multiple-choice formats (where evaluation involves comparing a predicted option such as “A/B/C/D” against a single-letter ground truth), real-world table tasks performed by professional experts are often more complex and nuanced. These tasks, such as code generation or structured reasoning, cannot be adequately evaluated using multiple-choice alone. In MMTU, we instead adopt a structured yet open-ended answer format (see Figure 1 for examples) for prediction and evaluation.
To support evaluations beyond simple string comparisons, we designed a lightweight evaluation framework that supports diverse evaluations in table tasks, including execution-based evaluation (for SQL and Python generation), and structured output evaluation (e.g., comparing an unordered JSON list against ground truth). Our framework is also extensible, making it easy to incorporate new task types and evaluation metrics. Details of our evaluation framework can be seen in [code-data].
@@ 4 Experiments
Experimental setup. In all our experiments reported below, we use publicly available model endpoints for inference [aoai-api, foundry-api] with default parameter settings. All of our code and data are publicly available at [code-data] for future research.
@@ 4.1 Overall performance
From a cost-efficiency perspective [^1] Price-per-token figures are sourced from the respective API providers [price-deepseek-r1, price-gpt-4o, price-gpt-4o-mini, price-llama3.1-8B, price-llama3.3-70B, price-mistral-large-2411, price-mistral-small-2503, price-o4-mini]., looking at the “cost per question” column in the table, several light-weight reasoning models (e.g., GPT-5-mini and Gemini-2.5-flash) are quite competitive in terms of both quality and cost (even after accounting for their intermediate thinking tokens), making them cost-effective choices for complex table tasks.
@@ 4.2 Detailed analysis and sensitivity
These findings suggest that, despite recent advances in long-context LLMs [long-context-2, Gpt-4-tr], long contexts remain a significant challenge for tables. In Appendix E, we use a detailed experimental comparison between the standard “needle-in-a-haystack (NIH)” [NIAH-orig, niah-2], and our table-based “needle-in-a-haystack in table (NIHT)” test – while frontier models perform nearly perfectly on NIH, their performance drops sharply on NIHT, revealing fundamental limitations in their ability to handle long table contexts.
Robustness to table permutation. Figure 7 illustrates model performance under different table permutations. Specifically, we randomly shuffle rows and/or columns in input tables [^2] In shuffling columns, we keep the 3 left-most columns in all tables intact, as these are likely the “key columns” or “entity columns” in tables [bing-assets, webtables] in tables, to best preserve the meaning of these tables., with the 4 bars in each group representing: (1) no shuffle, (2) row-only shuffle, (3) column-only shuffle, and (4) both row and column shuffle. Recall that unlike natural-language text, two dimension relational tables are permutation-invariant [table-gpt, cong2023observatory], meaning that shuffling rows and columns should generally not change their semantic meanings and the associated tasks (e.g., the tasks in Figure 1 will remain the same, even if the rows/columns in the associated tables are shuffled).
@@
Table format variations. Figure 8 shows different models’ performance using common table input formats: Markdown, CSV, JSON, and HTML. We observe that unlike prior studies that showed notable sensitivity of LLMs [table-llm] to table formats, our results indicate that today’s frontier models, especially reasoning ones (GPT-5 and R1), are becoming less sensitive to these format variations (except HTML that still lags behind other formats). This reduces the need for format-specific optimization as models continue to improve. For chat models (GPT-5-Chat and DeepSeek-V3), we note that using JSON has an advantage on MMTU, mainly because it is easier for models to identify value/column correspondence in the JSON format, especially in long table context settings (e.g., in needle-in-a-haystack tasks like NIHT and NIHI shown in Table 1).

### v4

@@ 1 Introduction
Remarkable progress has been made in foundation models [14, 15, 53, 121, 4], partly thanks to an expanding array of large-scale benchmarking efforts. Prominent examples include benchmarks for general language understanding (e.g., GLUE [124], Super-GLUE [123], BIG-Bench [115], MMLU [59], MMLU-pro [127]), as well as benchmarks focused on coding and STEM reasoning (e.g., SWE-bench [118], GPQA-diamond [106], AIME [3], LiveCodeBench [84]). These efforts have played a critical role in deepening our understanding and accelerating the progress of foundation models.
Tables and table-based use cases are central to many real-world applications, including spreadsheets [92, 48], databases [101, 114], and computational notebooks [68, 40], which often require expert-level users such as data engineers, analysts, and database administrators to operate. While LLMs have shown great promise in working with tables [46, 39, 47], existing evaluations of table-tasks remain narrow in scope – primarily focusing on NL-to-SQL [77, 147, 138] and Table-QA [137, 29, 24] – and fail to reflect the broader spectrum of real-world tasks that professional users face.
@@
Our evaluation shows that the complex and technical nature of MMTU demands a combination of capabilities – including table understanding, reasoning, and coding – that remain challenging for today’s frontier models. Even the top performing models, such as OpenAI GPT-5 and DeepSeek-R1 , achieve only 69.6% and 57.9% on MMTU, suggesting substantial room for improvement, and highlighting MMTU as a strong testbed for models aspiring toward general human-level intelligence, e.g., to meet or surpass the top 10% of skilled adults in diverse technical tasks like those characterized in [97].
@@
  • Reasoning models, such as OpenAI GPT-5 and DeepSeek R1, show a clear advantage over general-purpose chat models like GPT-5-Chat and DeepSeek-V3. The top reasoning models outperform the top chat models by over 10 percentage points (Table 3), underscoring the complexity of the tasks (which often require coding in SQL/Pandas) and the importance of reasoning in MMTU.
@@
Table Transform table-transform-by-relationalization (TTBR) Relationalize a table using transformation Acc [78, 10, 67, 132, 72] 230
table-transform-by-output-schema (TTBS) Synthesize transformation by output schema Acc [135, 110, 111] 685
table-transform-by-output-table (TTBT) Synthesize transformation by input/output tables Acc [11, 125, 122, 142] 86
Table Matching Entity matching (EM) Match rows refer to the same semantic entity Acc [38, 98, 146, 80, 102] 4228
Schema matching (SM) Match columns refer to the same concept F1 [71, 145, 12, 89] 669
Head value matching (HVM) Match column-headers with cell-values Acc [79, 105] 900
Data Cleaning data-imputation (DI) Predict missing values in tables Acc [16, 44, 79] 1971
error-detection (ED) Detect erroneous cells in tables F1 [31, 126, 90, 21] 1891
list-to-table (L2T) Split lists of undelimited values into table Acc [30, 17, 45, 75] 957
Table Join semantic-join (SJ) Predict semantic join between two tables Acc [57, 1, 37, 128] 130
equi-join-detect (EJ) Predict equi-joins between a set of tables F1 [81, 27, 108, 65] 494
Column Transform program-transform-by-example (PTBE) Program transformation by input/output examples Acc [56, 52, 43, 113, 58] 558
formula-by-context (FBC) Predict formula based on table context Acc [23, 25] 3513
semantic-transform-by-example (STBE) Predict semantic transformations by examples Acc [57, 1, 37] 131
Column Relationship arithmetic-relationship (AR) Predict arithmetic-relationship (AR) in tables F1 [112, 7] 818
functional-relationship (FR) Predict functional-relationship (FR) in tables F1 [7, 103, 129] 267
string-relationship (SR) Predict string-relationship (SR) in tables F1 [7, 56, 51] 765
Table understanding Needle-in-a-haystack-table (NIHT) Retrieve cell content in a table Acc [44] 999
Needle-in-a-haystack-index (NIHI) Retrieve index based on cell value Acc [44] 1000
NL-2-code NL-2-SQL (NS) Translate natural-language into SQL Acc [147, 77, 74, 138, 148] 2489
Table QA Table Question Answering (TQA) Answer questions based on tables Acc [131, 137, 28, 29] 1793
Fact Verification (FV) Verify facts based on tables Acc [24, 136, 141] 916
KB Mapping Column type annotation (CTA) Predict KB types based on column content Acc [66, 134, 116, 140, 61] 881
Column property annotation (CPA) Predict KB property for a pair of columns Acc [66, 36, 99] 873
Cell entity annotation (CEA) Predict KB entity for a table cell Acc [66, 42, 13, 130, 83] 892
@@ 2 Related Work
Large-scale benchmarks for foundation models. The rapid advancement of foundation models has made benchmark evaluation ever more important. Prominent large-scale benchmarks, such as GLUE [124] (9 NLP tasks), Super-GLUE [123] (10 NLP tasks), Big-BENCH [115] (204 tasks), MMMU [59] (15,908 questions), MMMU-pro [127] (12,032 questions), MMLU [139] (11,550 multi-modal questions) etc., offer comprehensive evaluations of model capabilities. However, as models improve rapidly, benchmarks can become saturated quickly (e.g., in the matter of a few years), prompting newer and more challenging benchmarks [123, 127]. All of these benchmarking efforts have nevertheless played a crucial role in measuring and stimulating the development of foundation models.
@@
Benchmarks for reasoning. Reasoning has recently emerged as a key challenge for foundation models. Beyond general intelligence benchmarks (e.g., GPQA Diamond [106], AGIEval [149], HLE [62]), there are specialized benchmarks targeting mathematical reasoning (e.g., AIME [3], MathVista [91], IMO [63]), coding (e.g., SWE-bench [118], CodeForce [34], LiveCodeBench [84], IOI [64]), and multi-modal reasoning (e.g., MMMU [139], ARC [5]). These benchmarks have become important tools for evaluating models’ ability to tackle complex reasoning tasks, but can also get saturated quickly (e.g., GSM8k [33], Math500 [60], HuamEval [20]), making it necessary to create new and more challenging benchmarks.
@@
Existing benchmarks relating to tables. Given the importance of table data, benchmarks have been developed in the ML and NLP community to evaluate model ability on tables, which however usually focus on a small set of table tasks such as NL-2-SQL [147, 77, 74, 138, 148] and Table-QA [24, 131, 137, 28, 29]. In contrast, MMTU expands the scope of current evaluations by incorporating 19 additional table tasks drawn from decades of research in communities such as data management and programming languages, resulting in a more comprehensive evaluation framework for assessing LLM capabilities on tabular data.
More recently, spreadsheet-centric benchmarks have emerged, including Spreadsheet-Bench [88], SheetCopilotBench [76], and Sheet-RM [26]. While these are important in the spreadsheet domain, these efforts are typically limited in scale (containing a few hundred cases), and are closely tied to specific file formats (e.g., .xlsx) and software environments. In comparison, MMTU focuses on general-purpose tabular data that applies broadly across spreadsheet, database, and computational notebook settings, enabling more scalable and format-independent evaluation of foundation models.
@@ 3.2 Data curation workflow
Literature survey. To ensure our table tasks reflect real-world challenges, we draw on our experience working on related problems, that many challenging predictive table tasks have been studied in the decades worth of computer science research, particularly in data management (SIGMOD/VLDB), programming languages (PLDI/POPL), and web data (WWW/WSDM) communities. We conduct a systematic survey of publications from these venues over the past two decades, leveraging a combination of keyword searches of paper titles, and DeepResearch-like tools [6] (where we specify detailed requirements for papers in these venues), to arrive at a promising set of papers and possible candidate tasks.
@@
  • (1) Real user-facing tasks, involving data tables that would otherwise require expert-level humans to perform. (We therefore exclude system-level predictive table tasks focused on performance improvements, such as query optimization [70, 73] and cardinality estimation [54, 69]);
  • (2) Objectively evaluable tasks, that come with unique manually-labeled ground truth. (We therefore exclude tasks such as table summarization [144, 9, 55, 104, 19] and table augmentation [143, 133], which lack unique ground-truth and may require subjective fuzzy LLM-based evaluations);
@@
To ensure the quality of MMTU, we conduct an LLM-based quality check using o4-mini. The model is prompted to evaluate each question for two aspects: (1) whether the question is ambiguous (since tasks like NL-code can sometimes be inherently ambiguous [109]), and (2) whether the ground-truth answer is correct. Approximately 8% of the original questions were flagged as either ambiguous or incorrect and subsequently removed to enhance benchmark quality.
@@
Evaluation framework. In contrast to benchmarks like MMLU [59] and MMMU [139], which primarily use multiple-choice formats (where evaluation involves comparing a predicted option such as “A/B/C/D” against a single-letter ground truth), real-world table tasks performed by professional experts are often more complex and nuanced. These tasks, such as code generation or structured reasoning, cannot be adequately evaluated using multiple-choice alone. In MMTU, we instead adopt a structured yet open-ended answer format (see Figure 1 for examples) for prediction and evaluation.
To support evaluations beyond simple string comparisons, we designed a lightweight evaluation framework that supports diverse evaluations in table tasks, including execution-based evaluation (for SQL and Python generation), and structured output evaluation (e.g., comparing an unordered JSON list against ground truth). Our framework is also extensible, making it easy to incorporate new task types and evaluation metrics. Details of our evaluation framework can be seen in [95].
@@ 4 Experiments
Experimental setup. In all our experiments reported below, we use publicly available model endpoints for inference [8, 96] with default parameter settings. All of our code and data are publicly available at [95] for future research.
@@ 4.1 Overall performance
From a cost-efficiency perspective [^1] Price-per-token figures are sourced from the respective API providers [41, 50, 49, 85, 86, 93, 94, 100]., looking at the “cost per question” column in the table, several light-weight reasoning models (e.g., GPT-5-mini and Gemini-2.5-flash) are quite competitive in terms of both quality and cost (even after accounting for their intermediate thinking tokens), making them cost-effective choices for complex table tasks.
@@ 4.2 Detailed analysis and sensitivity
These findings suggest that, despite recent advances in long-context LLMs [22, 2], long contexts remain a significant challenge for tables. In Appendix E, we use a detailed experimental comparison between the standard “needle-in-a-haystack (NIH)” [87, 107], and our table-based “needle-in-a-haystack in table (NIHT)” test – while frontier models perform nearly perfectly on NIH, their performance drops sharply on NIHT, revealing fundamental limitations in their ability to handle long table contexts.
Robustness to table permutation. Figure 7 illustrates model performance under different table permutations. Specifically, we randomly shuffle rows and/or columns in input tables [^2] In shuffling columns, we keep the 3 left-most columns in all tables intact, as these are likely the “key columns” or “entity columns” in tables [18, 16] in tables, to best preserve the meaning of these tables., with the 4 bars in each group representing: (1) no shuffle, (2) row-only shuffle, (3) column-only shuffle, and (4) both row and column shuffle. Recall that unlike natural-language text, two dimension relational tables are permutation-invariant [79, 35], meaning that shuffling rows and columns should generally not change their semantic meanings and the associated tasks (e.g., the tasks in Figure 1 will remain the same, even if the rows/columns in the associated tables are shuffled).
@@
Table format variations. Figure 8 shows different models’ performance using common table input formats: Markdown, CSV, JSON, and HTML. We observe that unlike prior studies that showed notable sensitivity of LLMs [117] to table formats, our results indicate that today’s frontier models, especially reasoning ones (GPT-5 and R1), are becoming less sensitive to these format variations (except HTML that still lags behind other formats). This reduces the need for format-specific optimization as models continue to improve. For chat models (GPT-5-Chat and DeepSeek-V3), we note that using JSON has an advantage on MMTU, mainly because it is easier for models to identify value/column correspondence in the JSON format, especially in long table context settings (e.g., in needle-in-a-haystack tasks like NIHT and NIHI shown in Table 1).